OpenStudy (anonymous):

Check my answer? Two similar cylinders have heights 6 inches and 15 inches. What is the ratio of their surface areas? A. 8 : 25 B. 9 : 10 C. 4 : 25 D. 9 : 25 I chose C.

OpenStudy (anonymous):

There's also another one. Gerry has two porcelain spheres. The radii of the spheres are in a ratio of 3 : 5. What is the ratio between the volumes of the two spheres? A. 27 : 125 B. 81 : 125 C. 3 : 5 D. 27 : 50 I chose B.

OpenStudy (anonymous):

Whoops, I mean A.

OpenStudy (anonymous):

Curved surface area, right?

OpenStudy (anonymous):

It doesn't say

OpenStudy (anonymous):

I don't think you can solve the first one without knowing the radius( be it CSA or TSA) But the second problem's answer is NOT Option B.

OpenStudy (anonymous):

The volume of a sphere is : \[\frac{ 4 }{ 3 }\pi r ^{3}\] Plug in the values and do it again.

OpenStudy (anonymous):

I did that and I got 113.04 : 523.3, and that's not one of the answers I must be doing something wrong.

ganeshie8 (ganeshie8):

both are right @Ber96

OpenStudy (anonymous):

Oh, ok. Thanks @ganeshie8

ganeshie8 (ganeshie8):

you dont wanto find volume here. just you wanto find the ratio of volumes. so you're original ticks are right.

OpenStudy (anonymous):

@ganeshie8 How's the second one correct? o_o

OpenStudy (anonymous):

I meant to say that my second answer was A, not B.

ganeshie8 (ganeshie8):

yep ! A is right for 2nd q. radii ratio = 3 : 5 volume ratio = 3^3 : 5^3

OpenStudy (anonymous):

Oh, sorry. I thought you said the answer was B. x-|

OpenStudy (anonymous):

It's alright, but thanks for your help, both of you :)

ganeshie8 (ganeshie8):

np :)